A series circuit has two resistor with resistances of 3,000Ω and 10,000Ω. If the current through the 3000 resistor is 1.5mA what is the voltage of the power supply

a. 4.00V b. 15.00V c. 3.46V d. 19.50V

There are 3 resistors hooked up in parallel. Their resistances are 500Ω 2,500Ω and 8,000Ω, if the power supply provides 10V, what is the current going through the 2500Ω resistor.

a.

0.9mA

b.

4.0mA

c.

1.2mA

d.

25.2mA

The power source of a circuit is 10V and the current is 2mA, what is the equivilent resistance in the circuit?

a.

.2 Ω

b.

5 Ω

c.

5000 Ω

d.

.2 mΩ
Answer #1

a.

from ohm's law,

V = I*Req

here, For series combination

Req = R1 + R2 + R3 +...............

given, R1 = 3000 ohm, R2 = 10000 ohm

then, Req = 13000 ohm

Remember in resistors parallel combination voltage distribution in each part will be same and in series combination current distribution in each resistor will be same. So in series circuit:

current in 3000 ohm = current in 10000 ohm = current in ciruit = 1.5 mA = 1.5*10^-3 A

So, V = (1.5*10^-3)*(13000)

V = 19.50 v

So, correct option is d.

b.)

As, in resistors parallel combination voltage distribution in each part will be same.

So, voltage across 500 ohm = voltage across 2500 ohm = voltage across 8000 ohm = power supply voltage = 10 V

Now, from ohm's law,

current in 2500 ohm(I) = (voltage across 2500 ohm)/resistance

I = 10/2500 = 0.004 A

I = 4.0 mA

So, correct option is b.

c.)

From ohm's law,

V = I*Req

here, V = power supply = 10 V

I = current in circuit = 2 mA = 2*10^-3 A

then, Req = equivalent resistance = V/I

Req = 10/(2*10^-3)

Req = 5000 ohm

So, correct option is c.

"Let me know if you have any query."